OpenStudy (anonymous):

Evaluate the expression, correct to six decimal places, by the Change of Base Formula and a calculator log2^3

jimthompson5910 (jim_thompson5910):

\[\Large \log_{2}(3) = \frac{\ln(3)}{\ln(2)}\] \[\Large \log_{2}(3) \approx \frac{1.098612}{0.693147}\] \[\Large \log_{2}(3) \approx 1.584962\]

OpenStudy (anonymous):

good job jim...you're really helping this precalc noob out

OpenStudy (anonymous):

never forget change of base \[\log_b(x)=\frac{\log_a(x)}{\log_a(b)}\] and since you only have \[\log_{10}\] and \[\ln\] on your calculator, those are the ones you can use. and don't forget what this really tells you. you just solved \[2^x=3\] for x via \[x=\frac{\ln(3)}{\ln(2)}\]